• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Roedelbronn

    Log entry time 00:03:15 on June29,2002

    Entry number 86994

    keyword=Swing shift summary

    16:00 Taking data for point RCS_C
    16:10 Stopped RCS data-taking to allow Dave Meekins to perform his target loop change.
    18:30 All work in the hall is done and we are going to Beam Permit. Loop 3 (bottom) is filled with H2, top has lead
    18:45 Hall is in Beam Permit
    18:58 We have beam; starting at 30uA with the 9-foil carbon optics target.
    19:01 Increasing current to 90uA
    19:10 Run # 1891................C optics run...............kinematics C_A
    19:25 Switching to cryo lead target, raster ON, 5 uA
    19:28 Run # 1892 ...............dummy run to check raster, raster size ~2 x 2.5
    19:31 Run # 1893................cryo lead tgt...............kinematics Pb_A
    19:56 Moved HRS-R to 60.559 degrees
    19:59 Run # 1894................cryo lead tgt...............kinematics Pb_B
    20:10 Current has been very steady at ~4.84 uA since the start of run 1892
    20:30 ROC 14 crash
    20:33 Run # 1895................cryo lead tgt...............kinematics Pb_B
    20:50 Moved HRS-R to 66.360 degrees
    20:52 Run # 1896................cryo lead tgt...............kinematics Pb_C
    21:28 Moved HRS-R to 72.160 degrees
    21:32 Run # 1897................cryo lead tgt...............kinematics Pb_D
    21:53 During run 1897, we are increasing current to 10 uA and then 15 uA.
    22:01 Back to 5 uA.
    22:30 Moved HRS-R to 60.559 degrees
    22:32 Run # 1898.........junk - prescalers not adjusted.
    22:35 Run # 1899................cryo lead tgt...............kinematics Pb_B
    22:58 Stut off beam in preparation for hall access. Access is to raise the trip point of the beam dump ion chamber dose rate from 5 rad/hr to 30 rad/hr
    23:00 Hall is in Controlled Access.
    23:36 Hall is back in Beam Permit, waiting for 25 uA.
    23:44 Run # 1900...............cryo lead tgt...............kinematics Pb_B